OpenStudy (anonymous):

bayside insurance offers two health palns under plan A Giselle would have to pay $150.00 of her medical bills plus 25% of the rest under plan B Gose;;e wpi;d [au tje first $190 but only 20% of the rest for what amount of medical bills will plan B save Giselle money? Assume she hasover $190 in bills.

OpenStudy (amistre64):

the trick here is not to read this to quickly

OpenStudy (amistre64):

plan A = (Bill - 150)(.25) + 150 plan B = (Bill - 190)(.20) + 190

OpenStudy (amistre64):

these can be rewritten in more of an xy nature

OpenStudy (amistre64):

y = (.25)x -150(.25) + 150 y = (.20)x -190(.20) + 190